-Equilibrium Concentration of [Fe(SCN)2+]   

– Reacted Concentration of Fe3+

– Reacted Concentration of SCN-

– Equilibrium Concentration of Fe3+

– Equilibrium Concentration of SCN-

– Equilibrium Constant, Kc

M(Fe) M(SCN)
0.00229 0.00206
Test Solutions Concentration Absorbance
1 0 0
2 0.0000618 0.154
3 0.000103 0.22
4 0.000144 0.352
5 0.000206 0.514

From Graph: y = 0.1226x - 0.1198
R² = 0.9844
Answer #1

Answer:

Given that your lab results:

From the graph of Absorbance (y-axis) versus concentration(x-axis) shown in attached excel image.

Plot Absorbance on y axis and concentration on x-axis.

You can draw trendline also.

The equation is y = 2483.23x = Beer's law constant

The slope is 2483.23

M(Fe3+) = 0.00229, M(SCN-) = 0.00206

The equilibrium reaction is

Fe3+ +. SCN- \rightleftharpoons .   FeSCN2+

I 0.00229 0.00206 0

C. -x. -x. +x

E. (0.00229-x) (0.00206-x). x

The expression for the equilibrium constant Kc is

Kc = [FeSCN2+]/ [Fe3+][SCN-]

Kc = x/(0.00229-x)(0.00206-x)............(1)

Here x is equilibrium concentration of FeSCN2+ which can be obtained by using Beer's law

A=\varepsilonlc

Where A is Absorbance

\varepsilonl = slope of the line = 2483.23

c = conc. of FeSCN2+

or c = A/slope = A/2483.23 .............(2)

since you have not given the value of A Absorptivity at which the equilibrium is reached. But if you have the value of A substitute in equation 2 to get the value of c or x.

x = A/2483.2318:18 * 44.1 VE 16 * 73 Book (8) Ą O O v v D. Office 365 feature - Get this feature and ... fx =SLOPE( B3:37 , A3:A7 ) А в 1

Equilibrium concentration of FeSCN2+ = x

Reacted concentration of fe3+ = SCN- = x

Equilibrium concentration of Fe3+ = (0.0029-x)

Equilibrium concentration of SCN- = (0.00206-x)

Substitute the values in equation 1 to get the value of Kc